OpenStudy (anonymous):

Which of the following numbers is irrational? A 5/9 B 3.15 C /5 D 8.341 I think its c please help

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thinks can you help me with some more please

OpenStudy (anonymous):

ig

OpenStudy (anonymous):

Which of the following expressions is equivalent to 2^{3}.2^{-9} A 1/64 B 1/32 C 1/16 D 1/8

OpenStudy (anonymous):

Which of the following represents the solution of the equation 2(x-4)=32? [8.EE.7.b] A. x=12 B. x=14 C. x=18 D. x=20

OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system of equations below? -3x+2y=2 7x-3y=7 [8.EE.8.b] A. (2, 4) B. (4, 7) C. (7, 14) D. (8, 13)